HUBBLE SPACE TELESCOPE DAILY REPORT #3083 PERIOD COVERED: 0000Z (UTC) 03/28/02 - 0000Z (UTC) 03/29/02 Daily Status Report as of 088/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.2 Completed Twelve Sets of NICMOS/1/2/3 8945 (Cooling System Monitoring) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3 was used 1}.to measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 2}. to demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature. and 3}.to demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. No problems were encountered. 1.3 Completed WF/PC-2 8950 (SM3B Cool Down, Contamination Monitor, and Focus Check) The WF/PC-2 was used to monitor the planned cooldown to -88C, followed by UV monitors, focus checks, as well as decons and associated monitors. No problems were noted. 1.4 Completed Three Sets of WF/PC-2 8698 (Identification of the Galaxy's Missing Mass) The WF/PC-2 was used to observe the nature of dark matter that is one of the key astrophysical questions of the day. The existence of dark matter and its dynamical dominance in the outer parts of our Galaxy and spiral galaxies with flat rotation curves is well established. The MACHO project has identified ~half of the Milky Way's dark matter with stellar objects of ~0.5 M_sun, probably white dwarfs. But the location of the microlensing in the halo is disputed. Several have detected two candidate halo white dwarfs of L/L_sun ~ 10^-5 in the Hubble Deep Field with 25 +/- 5 mas/year proper motions. The observations completed with no reported problems. 1.5 Completed Two Sets of STIS/CCD 8963 (CCD Bias) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD read noise from multiple bias frames obtained in common science modes: amplifier D, gains 1 and 4 electrons/DN, and binning factors of 1x1, 1x2, 2x1, and 2x2. The read noise measurements before and after NCC turn-on are then compared, and standard calibration reference files for use in pipeline reduction of SMOV3B data are constructed. The proposal had no reported problems. 1.6 Completed ACS/WFC/HRC 9031 (CCD Hot Pixel Annealing) The Advanced Camera for Surveys (WFC and HRC) was used to perform hot pixel annealing, which will be done once per month. The CCD TECs will be turned off and heaters will be activated to bring the WFC detector temperature to about +10C. The HRC temperature will reach about 30C. This state will be held for approximately 24 hours, after which the heaters are turned off, the TECs turned on, and the CCDs returned to normal operating condition. To assess the effectiveness of this procedure, a bias and two dark images will be taken before and after the annealing procedure for both WFC and HRC. There were no reported problems. 1.7 Completed STIS/CCD 8964 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain dark frames twice daily, as in a corresponding calibration program. Standard calibration reference darks are constructed and hot pixel lists are updated for use in pipeline reduction of SMOV3B data. All observations completed without incident. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 4 Successful: 4 Per HSTAR 8584, following a successful re-acquisition at 087/181114Z, a loss-of-lock occurred at 087/181352Z. No observations were affected. Scheduled Re-acquisitions: 10 Successful: 10 2.2 FHST Updates: Scheduled: 5 Successful: 5 2.3 Operations Notes: The NICMOS Cryo Cooler (NCC) cooldown continues nominally. SI C&DH errors were reset three times per ROP NS-5. The following Operations Notes were executed during this operational day; Adjust ACS Error Counter Limit @087/1446z Equipment Shelf by FHST Temp Limit Change @087/1448z NICMOS Safing Status Buffer Messages @087/1910z 3 Gyro Configurations Limit Changes @087/2057z Reset NICMOS MEB Radiator 2 Default Limit @087/2059z 3 Gyro Configurations Limit Changes @087/2343z Turn Off ACS HRC/SBC Corrector Status limit @088/0350z NICMOS Bad Mode Count OFF while NICMOS is safed @088/0359z Using ROP SR-1A, the SSR-1 EDAC error counter was cleared at 087/1013Z. The FHST 21 Shelf Temperature (ESHLF21T) flagged out-of-limits high at 087/122400Z with a value of 13.06189 degC (the limit is 13.0 degC). HSTAR 8583 was written. As directed by an operations request, the ACS Coarse Corrector Alignment -- Part 3 was performed at 087/1446Z. Per an operations request, Gyro 1 was powered off at 087/192058Z. Commanding to change the gyro configuration in the PSEA went smoothly, completing its reconfiguration at 087/192854Z. Final commanding to adjust gyro safemode test limits finished at 087/011418Z. Attitude errors from a Real Time Fine Attitude Determination showed approximately 11 arc-seconds of error (RSS) an hour after the gyro reconfiguration, well within the nominal errors. A second map showed slightly higher errors (~46 arc-seconds), not unexpected, and an ARU/PRT was scheduled to buy back additional margin to ensure the success of the Guide Star Acquisition. The following FHST Updates and Guide Star Acquisition were successful. SAC has fully reconfigured their software to match the current on-orbit configuration. These updated configuration files have been transported to all of the CCS operational strings. NICMOS was safed at 087/1910Z to help the NCS cooldown process. At 087/1910, using an operations request, STIS and ACS high voltage turn-ons (via SMS commanding, were inhibited until a sufficiently low temperature is reached during the NCC cooldown. A TTR was generated when a GCMR was not acknowledged by WSGT at 087/2302Z. ROP DF-18A was utilized to change an engineering status buffer limit System Monitor=11) at 088/0341Z. Per HSTAR 8585, the Zone 3B temperature (OTCZ3BT) began intermittently flagging out-of-limits low at 088/070758Z with a value of 17.8714 degC. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW